# The Identification Power of

### v1

@@ Abstract
This paper investigates the identification power gained by combining experimental data, in which treatment is randomized, with observational data, in which treatment is self-selected, for distributional treatment effect (DTE) parameters. While experimental data identify average treatment effects, many DTE parameters, such as the distribution of individual treatment effects, are only partially identified. We examine whether, and how, combining the two data sources tightens the identified set for these parameters. For broad classes of DTE parameters, we derive sharp bounds under the combined data and clarify the mechanism through which the data combination tightens the identified set relative to using experimental data alone. Our analysis highlights that self-selection in the observational data is a key source of identification power. We also characterize necessary and sufficient conditions under which the combined data shrink the identified set, showing that such shrinkage generally occurs unless selection-on-observables holds in the observational data. We also propose a linear programming approach to compute the sharp bounds, which can accommodate additional structural restrictions such as mutual stochastic monotonicity of potential outcomes and the generalized Roy model. An empirical application using data on negative campaign advertisements in a U.S. presidential election illustrates the practical relevance of the approach.
Keywords: Copula; Data combination; Heterogeneous treatment effect; Partial identification; Self-selection.
@@ 1 Introduction
In evaluating public policies and medical interventions, researchers often have access to both experimental and observational data. Experimental studies—such as randomized controlled trials (RCTs)—are widely regarded as the gold standard for causal inference, whereas observational data are far more prevalent in practice and often contain rich behavioral variation. This paper investigates the identification gains that arise from combining these two data sources, with a focus on distributional treatment effects.
@@
We consider a general setting in which treatment receipt and outcomes are observed from two data sources: experimental data, where treatment is randomly assigned, and observational data, where treatment is self-selected. This structure arises naturally in various policy and medical contexts. In development policy, for example, the efficacy of public health interventions such as mosquito nets is evaluated both through field experiments (where nets are randomly assigned) and through observational data (where nets are self-purchased). In education, randomized class-size assignment (e.g., Project STAR) coexists with parental or administrative decisions about classroom placements. In clinical medicine, drugs may be randomly assigned in clinical trials but, once approved, are prescribed based on physician judgment.
@@
Experimental data are valuable because the random assignment of treatment enables identification of key causal parameters, such as the average treatment effect (ATE). However, beyond the ATE, distributional treatment effect (DTE) parameters are often crucial to policy and medical evaluation, as well as to understanding treatment effect heterogeneity across individuals. Letting $Y_{1}$ and $Y_{0}$ denote potential outcomes under treatment and no-treatment, respectively, examples of such parameters include: (i) the fraction of individuals who benefit from treatment, $\mathbb{P}(Y_{1}>Y_{0})$; (ii) the distribution function of individual treatment effects, $F_{Y_{1}-Y_{0}}(\delta)\equiv\mathbb{P}(Y_{1}-Y_{0}\leq\delta)$; (iii) the ATE for disadvantaged individuals, $\mathbb{E}[Y_{1}-Y_{0}\mid Y_{0}\leq c]$, where the subpopulation with $Y_{0}\leq c$ represents individuals whose baseline outcomes fall below a threshold level $c$ (e.g., a poverty line); and (iv) the correlation between $Y_{1}$ and $Y_{0}$.
@@
Although these parameters are highly relevant for policy and medical evaluation, they are generally not point-identified from experimental data alone, as identification generally requires knowledge of the joint distribution of the two potential outcomes. Experimental data identify only the marginal distributions of $Y_{1}$ and $Y_{0}$, and thus yield partial identification (e.g., Heckman et al. (1997); Fan and Park (2010)). However, the resulting identified sets are often wide and, consequently, not very informative.
@@
This paper explores the extent to which combining experimental and observational data can improve identification for such distributional parameters. Specifically, we address the following questions: (i) Does combining the two data sources shrink the identified set for DTE parameters?; (ii) If so, what is the mechanism behind such shrinkage, and when does it occur?; (iii) How can we compute the identified set under such a data combination?
@@
To investigate these questions, we build on a framework that incorporates both experimental and observational data. In this setting, researchers observe individuals’ outcomes $Y$, treatment status $D\in\{0,1\}$, and a data source indicator $G\in\{\mathrm{exp},\mathrm{obs}\}$, denoting whether the observation comes from an experimental study ($G=\mathrm{exp}$) or an observational study ($G=\mathrm{obs}$). A key element of the framework is a latent self-selection variable $S\in\{0,1\}$, which represents the treatment choice an individual would make under self-selection. This variable is observed in the observational data (where $S=D$) but unobserved in the experimental data. We compare the identified sets under each data setting using a copula-based approach. Under standard assumptions—random treatment assignment in the experimental data and external validity across data sources— this framework enables a systematic comparison of what can be learned from experimental data alone versus from the combination of both data sources.
@@
Our central theoretical results show that the identified set under the combined data is generally smaller than that under the experimental data alone, except in special cases where the self-selection variable $S$ is independent of the potential outcomes. The key intuition is as follows. While experimental data identify only the marginal distributions $F_{Y_{1}}$ and $F_{Y_{0}}$ of the potential outcomes, combining the two data sources enables identification of the joint distributions $F_{Y_{1}S}$ and $F_{Y_{0}S}$ of $(Y_{1},S)$ and $(Y_{0},S)$, respectively. We also formally show that the distribution pairs $(F_{Y_{1}},F_{Y_{0}})$ and $(F_{Y_{1}S},F_{Y_{0}S})$ fully characterize the identified sets under the experimental and combined data, respectively. Because the self-selection variable $S$ encodes information about the dependence structure between the potential outcomes, knowledge of $(F_{Y_{1}S},F_{Y_{0}S})$ can tighten the set of feasible joint distributions of $(Y_{1},Y_{0})$.
@@
To quantify this identification gain and derive analytical expressions for the sharp bounds, we apply copula bound analysis by building on and extending Fan et al. (2017), who study the identifying power of covariates, to settings with a self-selection variable $S$. For broad classes of DTE parameters, represented by super-modular functions or $\varphi$-indicator functions (defined later), we derive explicit sharp bounds, which facilitate computing the identified set under data combination, and present necessary and sufficient conditions under which the data combination yields tighter identified sets. We show that combining the data sources improves identification when the dependence structure between the potential outcomes varies across self-selection types.
@@
To further tighten the identified sets, we also consider incorporating additional structural restrictions that are plausible in many empirical contexts. In particular, we consider two such restrictions: mutual stochastic monotonicity of the potential outcomes, as studied by Frandsen and Lefgren (2021), and a generalized Roy model condition. Though not required for our main results, these restrictions can substantially narrow the identified sets when imposed. To incorporate them under data combination, we develop a computational approach based on linear programming that efficiently computes sharp bounds for a broad class of DTE parameters.
@@
Our analysis also applies to data from Doubly Randomized Preference Trials (DRPTs) (Rücker, 1989), a unique yet recently prevalent design. In a DRPT, individuals are randomly assigned to one of three groups: a treatment group, a control (no-treatment) group, or a self-selection group, in which individuals choose between treatment and no-treatment. This design has been employed in a range of studies across the social sciences (Gaines and Kuklinski, 2011; Arceneaux et al., 2012; De Benedictis-Kessner et al., 2019; Ida et al., 2024) and medical sciences (King et al., 2005; Howard and Thornicroft, 2006). [^1] A known advantage of DRPTs is that they enable identification and estimation of the ATE for each self-selection group, $\mathbb{E}[Y_{1}-Y_{0}\mid S=s]$ for $s=0,1$; see, for example, Long et al. (2008) and Gaines and Kuklinski (2011).
@@
We illustrate the empirical relevance of our approach using the DRPT data from Gaines and Kuklinski (2011). That study examines the effects of negative campaign advertisements on individuals’ attitudes toward presidential candidates during the 2008 U.S. presidential election. The results show that incorporating self-selection data substantially narrows identified sets for DTE parameters such as $\mathbb{P}(Y_{1}<Y_{0})$, the fraction of individuals negatively affected by the advertisements. These findings highlight the practical value of combining random-assignment and self-selection data when estimating distributional treatment effects.
@@
This paper relates to and contributes to two strands of literature: (i) the combination of random-assignment and self-selection data sources for causal inference, and (ii) the partial or point identification of distributional treatment effects.
@@
The first strand of the literature explores the potential benefits of combining data from random-assignment and self-selection sources. This literature typically pursues two objectives: (1) improving the precision of estimates when experimental data suffice for identification, and (2) identifying causal effects that are not identifiable from experimental data alone. Regarding the first objective, Rosenman et al. (2023), Yang et al. (2023), Gui (2024), and Lin et al. (2025) propose novel estimation methods that improve the efficiency of treatment effect estimation through combining experimental and observational data. For the second objective, Long et al. (2008) show that the ATE for each self-selection group, $\mathbb{E}[Y_{1}-Y_{0}|S=s]$ for $s=0,1$, can be identified using data from a DRPT. Knox et al. (2019) extend this to multiple treatment settings and provide partial identification results. In a different context where experimental data contain secondary outcomes and observational data contain primary outcomes, Athey et al. (2025) show the identification of the ATE for the primary outcomes. Our setting differs from theirs, as we consider two data sources sharing the same outcome type.
@@
Our contribution to this strand of literature is to uncover a novel advantage of data combination: it can shrink the identified set for DTE parameters. In contrast to prior works focused on point identification of new parameters or improvements in estimation efficiency, we show that combining data improves informativeness by tightening bounds in partial identification settings.
@@
The second strand of the literature investigates the partial or point identification of distributional treatment effects under various assumptions and structural restrictions. [^2] Notable works include Heckman et al. (1997), Manski (1997), Fan and Park (2009; 2010; 2012), Fan and Wu (2010), Fan et al. (2014), Fan et al. (2017), Vuong and Xu (2017), Kim et al. (2018), Firpo and Ridder (2019), Mourifie et al. (2020), Callaway (2021), Frandsen and Lefgren (2021), Russell (2021), and Kaji and Cao (2023), among others. In particular, Fan and Park (2009; 2010; 2012), Fan et al. (2017), and Firpo and Ridder (2019) develop nonparametric bounds on the joint distribution of potential outcomes and its functionals under minimal assumptions, such as random treatment assignment. To tighten such bounds, subsequent works have explored additional structures, including structural models of self-selection (Fan and Wu, 2010; Mourifie et al., 2020), panel data with time-dependence restrictions (Callaway, 2021), mutual stochastic monotonicity of potential outcomes (Frandsen and Lefgren, 2021), and so on.
@@
This paper contributes to this literature by clarifying the identifying power that arises from combining experimental and observational data—a novel structure not previously explored—and by providing a computable characterization of the sharp bounds under such data combination.
@@
The remainder of the paper is organized as follows. Section 2 introduces the data combination framework, formalizes the DTE parameters, and define their identified sets under the experimental data and the combined data, respectively. Section 3 characterizes the identified sets for each data scenario, highlighting the sources of identification gains. Section 4 derives sharp bounds for broad classes of DTE parameters—those represented by super-modular or $\varphi$-indicator functions— and establishes necessary and sufficient conditions under which data combination improves identification. Section 5 introduces additional restrictions and develops a linear programming approach to compute sharp bounds. Section 6 presents an empirical application using DRPT data from Gaines and Kuklinski (2011). Section 7 concludes. All proofs are provided in the appendix.
@@ 2.1 Combined Experimental and Observational Data
We suppose that the observed data consist of the quadruple $(Y,D,G,X)$, where $Y$ is the observed outcome; $D\in\{0,1\}$ is a binary treatment indicator; $G\in\{\mathrm{exp},\mathrm{obs}\}$ indicates the data source; and $X$ denotes a vector of covariates (pre-treatment variables). Specifically, $G$ indicates whether each observation originates from an experimental study ($G=\mathrm{exp}$) or an observational study ($G=\mathrm{obs}$). The treatment indicator $D$ equals 1 if the individual receives the treatment and 0 otherwise. In the experimental data, $D$ is randomly assigned, whereas in the observational data, it is determined by self-selection. Let $Y_{1}$ and $Y_{0}$ be continuous random variables representing the potential outcomes under treatment and no-treatment, respectively. The observed outcome is defined as $Y\equiv DY_{1}+(1-D)Y_{0}$. We denote the support of $X$ by $\mathcal{X}$.
We introduce a latent self-selection variable $S\in\{0,1\}$ that represents the treatment choice an individual would make if he or she self-selects. In the observational data ($G=\mathrm{obs}$), we suppose that $S$ coincides with the observed treatment, i.e., $S=D$. In the experimental data ($G=\mathrm{exp}$), $S$ is unobserved. The variable $S$ can also be interpreted as a latent preference for treatment versus no-treatment.
Let $F^{\ast}$ denote the true cumulative distribution function (CDF) of all defined variables $(Y_{1},Y_{0},Y,D,S,G,X)$, including both observed and unobserved ones. For any generic CDF $F$, we denote probability and expectation under $F$ by $\mathbb{P}_{F}(\cdot)$ and $\mathbb{E}_{F}[\cdot]$, respectively. Specifically, for the true CDF $F^{\ast}$, we use the shorthand $\mathbb{P}(\cdot)$ and $\mathbb{E}[\cdot]$ in place of $\mathbb{P}_{F^{\ast}}(\cdot)$ and $\mathbb{E}_{F^{\ast}}[\cdot]$.
Throughout the paper, we suppose that the true CDF $F^{\ast}$ satisfies the following assumptions.
Assumption 2.1 (Potential Outcomes).
@@
Assumption 2.2 states that self-selection corresponds to the received treatment in the observational data, which is trivially satisfied by the definition of $S$. Assumption 2.3 requires that treatment is randomly assigned in the experimental study, possibly conditional on the pre-treatment variables $X$. Assumption 2.4 concerns the external validity of the data sources, requiring that systematic difference between the populations in the experimental and observational studies are captured by the pre-treatment variables $X$. This assumption is automatically satisfied under a DRPT design, where the data source indicator $G$ is randomly assigned; specifically, individuals are randomly assigned to the treatment, no-treatment, and self-selection groups. Assumption 2.5 imposes an overlap condition on both the treatment status and data source. [^3] This overlap condition can be relaxed at the cost of obtaining a wider identified set from the combined data. For instance, if a covariate value $x$ does not satisfy $\mathbb{P}(D=d,G=g\mid X=x)>0$ for some $d$ or $g$, then experimental and observational data cannot be combined for that value of $x$. However, one may still use either the experimental or the observational data alone to construct an identified set conditional on $x$.

### v4

@@ Abstract
This study investigates the identification power gained by combining experimental data, in which treatment is randomized, with observational data, in which treatment is self-selected, for distributional treatment effect (DTE) parameters. While experimental data identify average treatment effects, many DTE parameters—such as the distribution of individual treatment effects—are only partially identified. We examine whether and how combining these two data sources tightens the identified set for such parameters. For broad classes of DTE parameters, we derive nonparametric sharp bounds under the combined data and clarify the mechanism through which data combination improves identification relative to using experimental data alone. Our analysis highlights that self-selection in observational data is a key source of identification power. We establish necessary and sufficient conditions under which the combined data shrink the identified set, showing that such shrinkage generally occurs unless selection-on-observables holds in the observational data. We also propose a linear programming approach to compute sharp bounds that can incorporate additional structural restrictions, such as positive dependence between potential outcomes and the generalized Roy selection model. An empirical application using data on negative campaign advertisements in the 2008 U.S. presidential election illustrates the practical relevance of the proposed approach.
Keywords: Data combination; Distributional treatment effect; Heterogeneous treatment effect; Partial identification; Self-selection.
@@ 1 Introduction
Researchers often have access to both experimental and observational data when evaluating public policies and medical interventions. Experimental studies, such as randomized controlled trials, are regarded as the gold standard for causal inference, whereas observational data are far more prevalent in practice and often contain rich behavioral variation. This study investigates the identification gains that result from combining these two data sources, with a focus on distributional treatment effects (DTEs).
@@
We consider a general setting in which treatment receipt and outcomes are observed from two data sources: experimental data, where the treatment is randomly assigned, and observational data, where the treatment is self-selected. This structure arises naturally in various policy and medical contexts. In development policy, for example, the efficacy of public health interventions such as mosquito nets is evaluated through field experiments (where nets are randomly assigned) and observational data (where nets are self-purchased). In education, randomized class-size assignments (e.g., Project STAR) coexist with parental or administrative decisions regarding classroom placement. In clinical medicine, drugs may be randomly assigned in clinical trials; however, once approved, they are prescribed based on a physician’s judgment.
@@
Experimental data are valuable because the random assignment of treatment enables identification of key causal parameters such as the average treatment effect (ATE). However, beyond the ATE, DTE parameters are often crucial to policy and medical evaluation, as well as for understanding treatment effect heterogeneity. Let $Y_{1}$ and $Y_{0}$ denote potential outcomes under treatment and no-treatment, respectively. Examples of such parameters include (i) the fraction of individuals who benefit from treatment $\mathbb{P}(Y_{1}>Y_{0})$; (ii) distribution function of individual treatment effects $F_{Y_{1}-Y_{0}}(\delta)\equiv\mathbb{P}(Y_{1}-Y_{0}\leq\delta)$; (iii) ATE for disadvantaged individuals $\mathbb{E}[Y_{1}-Y_{0}\mid Y_{0}\leq c]$, where the subpopulation with $Y_{0}\leq c$ represents individuals whose baseline outcomes fall below a threshold level $c$ (e.g., a poverty line); and (iv) the correlation between $Y_{1}$ and $Y_{0}$.
@@
Although these parameters are highly relevant for policy and medical evaluations, they are generally not point-identified from experimental data alone because their identification requires knowledge of the joint distribution of the two potential outcomes. Experimental data identify only the marginal distributions of $Y_{1}$ and $Y_{0}$ and hence yield partial identification (e.g., Heckman et al. (1997); Fan and Park (2010)). However, the resulting identified sets are often wide and not very informative.
@@
This study explores whether and how combining experimental and observational data can improve the identification of distributional parameters. Specifically, we address the following questions. (i) Does combining the two data sources shrink the identified set for DTE parameters? (ii) If so, what mechanism drives this shrinkage, and under what conditions does it occur? (iii) How can the identified set under the combined data be characterized and computed?
@@
To address these questions, we build on a framework that incorporates both experimental and observational data. In this setting, researchers observe individuals’ outcomes $Y$, treatment status $D\in\{0,1\}$, and a data source indicator $G\in\{\mathrm{exp},\mathrm{obs}\}$, which indicates whether an observation comes from an experimental study ($G=\mathrm{exp}$) or an observational study ($G=\mathrm{obs}$). A key element of the framework is the latent self-selection type $S\in\{0,1\}$, which represents the treatment choice an individual would make under self-selection. This variable is observed in the observational data (where $S=D$) but unobserved in the experimental data.
We define the identified sets under experimental data alone and under the combined data, and then use a copula-based approach to characterize and compare them. Under standard assumptions—random treatment assignment in the experimental data and external validity across data sources—this framework enables a systematic comparison of what can be learned from experimental data alone versus combining both data sources.
@@
Our central theoretical results show that the identified set under the combined data is generally smaller than that under the experimental data alone, except in special cases where the latent self-selection $S$ is independent of the potential outcomes. The intuition is as follows. While the experimental data identify only the marginal distributions, $F_{Y_{1}}$ and $F_{Y_{0}}$, of the potential outcomes, combining the two data sources enables identification of the joint distributions $(F_{Y_{1}S},F_{Y_{0}S})$ of each potential outcome and the self-selection type. We show that the distribution pairs $(F_{Y_{1}},F_{Y_{0}})$ and $(F_{Y_{1}S},F_{Y_{0}S})$ fully characterize the identified sets under the experimental and combined data, respectively. As the latent self-selection $S$ encodes information about the dependence structure between the potential outcomes, knowledge of $(F_{Y_{1}S},F_{Y_{0}S})$ can tighten the set of feasible joint distributions of $(Y_{1},Y_{0})$.
@@
To quantify this identification gain and derive nonparametric sharp bounds, we apply copula bound analysis that builds on and extends the work of Fan et al. (2017), who study the identifying power of covariates, to settings with the self-selection variable $S$. For broad classes of DTE parameters represented by supermodular functions or $\varphi$-indicator functions (defined later), we derive analytical expressions for the sharp bounds and establish necessary and sufficient conditions under which the data combination yields tighter identified sets. We show that combining the two data sources improves identification when the dependence structure between the potential outcomes varies across the latent self-selection types.
@@
To further tighten the identified sets, we also consider incorporating additional structural restrictions that are plausible in many empirical contexts. In particular, we consider two such restrictions: positive dependence between the potential outcomes (Joe, 2014; Frandsen and Lefgren, 2021) and a generalized Roy selection model. Though not required for our main results, these restrictions can substantially narrow the identified sets. To implement these restrictions under the data combination, we develop a linear programming approach that efficiently computes sharp bounds for a broad class of DTE parameters.
@@
Our analysis also extends to data from doubly randomized preference trials (DRPTs) (Rücker, 1989; Long et al., 2008), a unique yet recently prevalent design. In DRPTs, individuals are randomly assigned to one of three groups: a treatment group, a control (no-treatment) group, or a self-selection group, in which individuals choose between treatment and no-treatment. This design has been used in a wide range of studies across the social sciences (Gaines and Kuklinski, 2011; Arceneaux et al., 2012; De Benedictis-Kessner et al., 2019; Ida et al., 2025) and medical sciences (King et al., 2005; Howard and Thornicroft, 2006). An established advantage of DRPTs is that they enable identification of the ATE for each self-selection group, $\mathbb{E}[Y_{1}-Y_{0}\mid S=s]$ for $s\in\{0,1\}$ (Long et al., 2008). Our results highlight a novel advantage of this design: it improves the identification of DTEs by combining random-assignment and self-selection samples.
@@
We illustrate the empirical relevance of our approach using DRPT data from Gaines and Kuklinski (2011), who study the effects of negative campaign advertisements on individuals’ attitudes toward presidential candidates during the 2008 U.S. presidential election. We find that incorporating self-selection data substantially tightens the identified sets for DTE parameters such as $\mathbb{P}(Y_{1}<Y_{0})$, the fraction of individuals negatively affected by the advertisements. The results further indicate that the advertisements have substantial but heterogeneous effects. These findings highlight the empirical value of combining random-assignment and self-selection data for estimating DTEs.
@@
This study relates and contributes to two strands of literature: (i) the combination of random-assignment and self-selection data sources for causal inference, and (ii) partial or point identification of DTEs.
@@
The first strand explores the benefits of combining data from random-assignment and self-selection sources. This literature typically pursues two objectives: (1) improving the precision of estimates when experimental data suffice for identification and (2) identifying causal effects that are not identifiable from experimental data alone. Regarding the first objective, Rosenman et al. (2023), Yang et al. (2023), and Gui (2024) propose novel estimation methods that improve the efficiency of treatment effect estimation by combining experimental and observational data. Regarding the second objective, Long et al. (2008) show that the ATE for each self-selection group, $\mathbb{E}[Y_{1}-Y_{0}|S=s]$ ($s=0,1$), can be identified using data from a DRPT. Knox et al. (2019) extend this to multiple treatment settings and derive partial identification results. In a different context, where experimental data contain secondary outcomes and observational data contain primary outcomes, Athey et al. (2025) study identification of the ATE for the primary outcomes. Our setting differs from theirs in that we consider two data sources sharing the same outcome type.
@@
Our contribution to this strand of the literature is to uncover a novel advantage of data combination: it can shrink the identified sets for DTE parameters. Unlike prior studies that focus on point identification of new parameters or gains in estimation efficiency, we show that combining data improves informativeness by tightening bounds in partially identified settings.
@@
The second strand studies the partial or point identification of DTEs under various assumptions and structural restrictions. [^1] Notable works include Heckman et al. (1997), Manski (1997), Fan and Park (2009; 2010; 2012), Fan and Wu (2010), Fan et al. (2014), Fan et al. (2017), Vuong and Xu (2017), Kim et al. (2018), Firpo and Ridder (2019), Mourifie et al. (2020), Callaway (2021), Frandsen and Lefgren (2021), Russell (2021), Lee (2024), Cui and Han (2025), and Kaji and Cao (2025) among others. In particular, Fan and Park (2009; 2010; 2012), Fan et al. (2017), and Firpo and Ridder (2019) develop nonparametric bounds for the joint distribution of potential outcomes and its functionals under minimal assumptions such as random treatment assignment. To tighten these bounds, subsequent studies have explored additional structures, including self-selection models (Fan and Wu, 2010; Mourifie et al., 2020), panel data with time-dependence restrictions (Callaway, 2021), mutual stochastic monotonicity of potential outcomes (Frandsen and Lefgren, 2021), and so on.
@@
This study contributes to this literature by clarifying the identifying power that arises from combining experimental and observational data—a novel structure not previously explored—and by providing computable characterizations of the sharp bounds under the data combination.
Finally, this study contributes to the broader causal inference literature by offering a new perspective on the role of observational data in identification. While the prevailing view holds that observational data become redundant for identification once experimental data are available, this study shows that they can nonetheless enhance the identification of DTEs, even in the presence of experimental data.
@@
The remainder of the paper is organized as follows. Section 2 introduces the data combination framework, formalizes the DTE parameters, and defines their identified sets under experimental data alone and under the combined data. Section 3 characterizes the identified sets under each data scenario and highlights the sources of identification gains from data combination. Section 4 derives sharp bounds for broad classes of DTE parameters, specifically those represented by super-modular or $\varphi$-indicator functions, and establishes necessary and sufficient conditions under which data combination improves identification. Section 5 introduces additional restrictions and develops a linear programming approach to compute sharp bounds. Section 6 provides an empirical illustration using DRPT data from Gaines and Kuklinski (2011). Section 7 concludes. All proofs and numerical examples are provided in the appendix.
@@ 2 Setup
We begin by outlining the framework. Section 2.1 introduces the data combination setting and fundamental assumptions. Section 2.2 defines the DTE parameters and provides illustrative examples. Section 2.3 defines the identified sets under experimental and combined data.
@@ 2.1 Combined Experimental and Observational Data
We consider observed data consisting of the quadruple $(Y,D,G,X)$, where $Y$ is the outcome; $D\in\{0,1\}$ is a binary treatment indicator; $G\in\{\mathrm{exp},\mathrm{obs}\}$ denotes the data source; and $X$ is a vector of pre-treatment covariates with its support denoted by $\mathcal{X}$. Specifically, $G$ indicates whether an observation comes from an experimental study ($G=\mathrm{exp}$) or observational study ($G=\mathrm{obs}$). The treatment indicator $D$ equals $1$ if the individual receives treatment and 0 otherwise. In the experimental data ($G=\mathrm{exp}$), $D$ is randomly assigned, whereas in the observational data ($G=\mathrm{obs}$), it is determined through self-selection. Let $Y_{1}$ and $Y_{0}$ denote the potential outcomes under treatment and no-treatment, respectively, both of which are assumed to be continuous. The observed outcome is defined as $Y\equiv DY_{1}+(1-D)Y_{0}$.
We introduce a latent self-selection variable $S\in\{0,1\}$ that represents the treatment choice an individual would make if he or she self-selects. In the observational data ($G=\mathrm{obs}$), $S$ coincides with the observed treatment, i.e., $S=D$, whereas in the experimental data ($G=\mathrm{exp}$), $S$ is unobserved. The variable $S$ can also be interpreted as a latent preference for treatment versus no-treatment.
Let $F^{\ast}$ denote the true cumulative distribution function (CDF) of all defined variables $(Y_{1},Y_{0},Y,D,S,G,X)$, including both observed and unobserved ones. For a generic CDF $F$, we denote the probability and expectation under $F$ by $\mathbb{P}_{F}(\cdot)$ and $\mathbb{E}_{F}[\cdot]$, respectively. For the true CDF $F^{\ast}$, we use the shorthand $\mathbb{P}(\cdot)$ and $\mathbb{E}[\cdot]$ in place of $\mathbb{P}_{F^{\ast}}(\cdot)$ and $\mathbb{E}_{F^{\ast}}[\cdot]$.
Throughout the paper, we suppose that $F^{\ast}$ satisfies the following assumptions.
Assumption 2.1 (Oberved Outcomes).
@@
Assumption 2.2 states that self-selection corresponds to the received treatment in the observational data, which is trivially satisfied by the definition of $S$. Assumption 2.3 requires that treatment is randomly assigned in the experimental study, possibly conditional on $X$. Assumption 2.4 concerns the external validity of the data sources, requiring that any systematic difference between the populations in the experimental and observational studies are captured by $X$. This assumption is automatically satisfied under a DRPT design, where the data source indicator $G$ is randomly assigned. In such a design, individuals are randomly assigned to treatment, no-treatment, and self-selection groups. Assumption 2.5 imposes an overlap condition on both the treatment status and data source. [^2] This overlap condition can be relaxed at the cost of yielding wider identified sets under the combined data. For example, if a covariate value $x$ does not satisfy $\mathbb{P}(D=d,G=g\mid X=x)>0$ for some $g$, then experimental and observational data cannot be combined at that value of $x$. In this case, one may still rely on either the experimental or the observational data alone to construct an identified set conditional on $x$.